Find the Perimeter of the Parallelogram

Top & Bottom:20.2/Sides:11.4

Found 2 solutions by RAY100, rfer:
Answer by RAY100(1637)

Perimeter = 2l +2w = 2(20.2) + 2(11.4) = 40.4 + 22.8 = 63.2
.

Answer by rfer(16322)

p=2(20.2)+2(11.4)=63.2